# Colors of Life in the Clouds: Biopigments of atmospheric microorganisms as a new signature to detect life on planets like Earth

### v2

@@ Appendix A Information on the strains collected in the atmosphere
L9-4, a closely related strain to Modestobacter versicolor (98.7% 16S rRNA identity), a microorganism that, as the name suggests, produces different pigments, namely both pink/red carotenoids, and melanin (black) for protection against radiation and lack of resources (G. S. N. Reddy et al., 2007). L9-9A-1 is closely related to Roseomonas vinacea (Y.-Q. Zhang et al., 2008) (98.0% 16S rRNA identity), a dark pink species. L9-7, a strain closely related to Micrococcus luteus (C.-H. Huang et al., 2019) (98.3% 16S rRNA identity), which, as the name indicates, is yellow. L6-1 is Curtobacterium aetherium (J. Mijatović Scouten et al., 2025), and L7-5 Curtobacterium oceanosedimentum (98.3% 16S rRNA identity) are yellow/orange bacteria extremely resilient (G. S. N. Reddy et al., 2007) to desiccation and UV-C radiation in the vegetative state, even more so than endospores (N. Bryan, 2017; P. Evseev et al., 2022). L7-1 is closely related to Massilia niabensis (99.2% 16S rRNA identity), the only specimens of this group that have been initially isolated from aerosols. Lastly, L7-7A is closely related to the yellow Noviherbaspirillum soli (S.-Y. Lin et al., 2013) (99.6% 16S rRNA identity).

### v3

@@ Appendix A Information on the strains collected in the atmosphere
L9-4 is a a closely related strain to Modestobacter versicolor (98.7% 16S rRNA identity), a microorganism that, as the name suggests, produces different pigments, namely both pink/red carotenoids, and melanin (black) for protection against radiation and lack of resources (G. S. N. Reddy et al., 2007). L9-9A-1 is closely related to Roseomonas vinacea (Y.-Q. Zhang et al., 2008) (98.0% 16S rRNA identity), a dark pink species. L9-7, a strain closely related to Micrococcus luteus (C.-H. Huang et al., 2019) (98.3% 16S rRNA identity), which, as the name indicates, is yellow. L6-1 Curtobacterium aetherium (J. Mijatović Scouten et al., 2025), and L7-5 Curtobacterium oceanosedimentum (98.3% 16S rRNA identity) are yellow/orange bacteria extremely resilient (G. S. N. Reddy et al., 2007) to desiccation and UV-C radiation in the vegetative state, even more so than endospores (N. Bryan, 2017; P. Evseev et al., 2022). L7-1 is closely related to Massilia niabensis (99.2% 16S rRNA identity), the only specimens of this group that have been initially isolated from aerosols. Lastly, L7-7A is closely related to the yellow Noviherbaspirillum soli (S.-Y. Lin et al., 2013) (99.6% 16S rRNA identity).